OpenStudy (anonymous):

How much pure acid should be mixed with 6 gallons of a 60% acid solution in order to get an 80% acid solution?

OpenStudy (anonymous):

I really need help!!

ganeshie8 (ganeshie8):

To start with, you have 6 gallons of a 60% acid solution. Whats 60% of 6 gallons ?

OpenStudy (anonymous):

3.6?

ganeshie8 (ganeshie8):

Yes, so we have 3.6 gallons of acid in the given 6 gallon solution

ganeshie8 (ganeshie8):

Lets pour some pure acid and increase the concentration to 80%

ganeshie8 (ganeshie8):

As usual, let \(x\) be the amount of extra acid that we need to pour to increase the concentration to 80%

ganeshie8 (ganeshie8):

Then in the end, we will be having \(x+3.6\) gallons of pure acid in the \(x+6\) gallon solution, yes ?

ganeshie8 (ganeshie8):

We want the acid in this final solution to be 80% : \[x+3.6 = \text{80% of }(x+6)\]

ganeshie8 (ganeshie8):

\[x+3.6 = 0.8(x+6)\] you can solve \(x\)

OpenStudy (anonymous):

Okay, wait, i'm about to do it.

ganeshie8 (ganeshie8):

Okay, take your time

OpenStudy (anonymous):

I got 6.

ganeshie8 (ganeshie8):

Correct!

OpenStudy (anonymous):

Thanks a lot!!

ganeshie8 (ganeshie8):

So, the final solution will have 6+3.6 = 9.6 gallons of acid And the solution would be 6+6 = 12 gallons

ganeshie8 (ganeshie8):

To verify your answer, you may simply find how much percent 9.6 is of 12

ganeshie8 (ganeshie8):

you should get 80%